A truck carrying 3600 cubic feet of cargo consisting of washing machines and refrigerators was hijacked. The washing machine are worth $300 each and are shipped in 36 cubic-foot cartons. The refrigerators are worth $900 each and are shipped in 45 cubic-foot cartons. If the total value of the cargo is $51,000, then how many of each were there on the truck?
Answer by stanbon(75887)

A truck carrying 3600 cubic feet of cargo consisting of washing machines and refrigerators was hijacked. The washing machine are worth $300 each and are shipped in 36 cubic-foot cartons. The refrigerators are worth $900 each and are shipped in 45 cubic-foot cartons. If the total value of the cargo is $51,000, then how many of each were there on the truck?
-----------------
value equation: 300w + 900r = 51000
volume equation: 36w + 45r = 3600
=======================================
Rearrange:
w + 3r = 170
4w + 5r = 400
-------------------
Multiply thru 1st by 4 to get:
4w + 12r = 4*170
4w + 5r = 400
--------------------
Subtract 2nd from 1st to solve from r:
7r = 4*170-400
r = 40 (# of refrigerators)
w = 170-3r = 170-3*40 = 50 (# of washers)
===================================
Cheers,
Stan H.